LOG-IN
Displaying reports 61-80 of 3360.Go to page 1 2 3 4 5 6 7 8 9 10 End
BIGFOOT (Readout)
Print this report.
ShalikaSingh - 15:28, Thursday 06 February 2025 (3901)Get code to link to this report
Telescope for EOM

Considering the beam after the 1st telescope (after the FI), with characteristics of w0=403.3um at 792.4mm after the telsecope --> Fig 1. 

We need to fit 2 EOM with aperture 2mm diameter, and around 100mm width. 

For 2 EOM, this space is 200 mm. Considering the space to mount them 250 mm is reserved

With a polarizer and the photodiode, the total space should be around 350mm.

The photodiode has aperture of 500 um diameter.

The beam should be hence, around 500/2/2 um = 125 um, radius. It was not possible to have lower beam waist with available lens. A lens of -200 mm focal length can perhaps help us do this.

The telescope was designed using Jammt. The Fig 2 shows the expected beam characteristics.

The first lens F=-100mm, is kept at roughly 140 mm after the initial telescope. The 2nd lens F=100mm is kept at 180 mm. There is a label on table after the 1st telescope defining the 0th position .

After that QWP and HWP are installed to make linear polarization. They are at 200-220 mm.

The EOM is kept at 400-500 mm at 45degree for characterization.

The beam characteristics after the 2nd telescope is shown, in Fig 2.

Images attached to this report
3901_20250204060051_modifiedtelescope.png 3901_20250206072836_beam.png
BIGFOOT (General)
Print this report.
MarcEisenmann - 23:50, Wednesday 05 February 2025 (3903)Get code to link to this report
Comment to SLM installation (Click here to view original report: 3899)

In order to properly connect to the SLM through the DVI cable we somehow need the usb controller vi.

It allows to select which output we want to use (usb/dvi).

If connection is properly made, the vi can read the SLM and controller temperature.

KAGRA MIR (General)
Print this report.
MarcEisenmann - 23:46, Wednesday 05 February 2025 (3902)Get code to link to this report
restart of pci lab

As we have transformed the previous TAMA storage into a now operational lab space we can now restart the PCI activities.

I've reconnected all required cables, restarted lockin, chopper and laser controllers.

Somehow, the COM port of the translation stage changed from COM6 to COM7.

I've re-set up it on this new COM port (baud rate 9600) with same settings as previously used.

I home the translation stage after moving it by hand outside of the optical components.

BIGFOOT (Readout)
Print this report.
ShalikaSingh - 17:46, Monday 03 February 2025 (3900)Get code to link to this report
EOM characterization with modified circuit

The impedance matching with EOM is now done using a 50ohm transformer, T-626-X65+. I am not using the opamp anymore, since we can't use upto 10Vp-p from moku (amplitude 5V).

The L = 13.6mH, and with EOM around 12pF, the resonant frequency for ideal circuit is around 0.39MHz. 

The circuit was supplied 10vp-p with moku and the spectrum of PD was read using moku. From the measurement it seems the resonant frequency is around 0.335MHz.

Fig 1: Spectrum of PD with laser off, and eom voltage off

Fig 2: Specturm of PD with laser on, eom voltage off

Fig 3: Spectrum of PD with laser on, eom voltage on.

Zoomed around the resonant freq to have better resolution of the peak. Then averaged 20 frames. 

Fig 4: Spectrum of PD with laser off, and eom voltage off

Fig 5: Specturm of PD with laser on, eom voltage off

Fig 6: Spectrum of PD with laser on, eom voltage on.

Images attached to this report
3900_20250203094503_pdspectrumlaseroff.png 3900_20250203094508_pdspectrumlaseroneomoff.png 3900_20250203094512_pdspectrumlaseroneomon.png 3900_20250203101704_pdspectrumlaseroffzoomed.png 3900_20250203101803_pdspectrumlaseroneomoffzoomed.png 3900_20250203101808_pdspectrumlaseroneomonzoomed.png
BIGFOOT (General)
Print this report.
MarcEisenmann - 23:30, Thursday 30 January 2025 (3899)Get code to link to this report
SLM installation

[Frederic, Marc]

The SLM came with a CD containing installation software, manual and correction masks.

They are saved in dropbox and NAS in 'SLM' folder.

Note that the correct installation file is not the x64 one.

Comments related to this report
MarcEisenmann - 23:50, Wednesday 05 February 2025 (3903)

In order to properly connect to the SLM through the DVI cable we somehow need the usb controller vi.

It allows to select which output we want to use (usb/dvi).

If connection is properly made, the vi can read the SLM and controller temperature.

BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 18:56, Thursday 30 January 2025 (3897)Get code to link to this report
hydrogel di-phenyl-di-peg

Katsuki, Yanada, Taki, Shalika

di-phenyl-di-peg(YI044)

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\Thu, Jan 30, 2025 4-51-15 PM.txt

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\Thu, Jan 30, 2025 5-18-49 PM.txt

input measument:

C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\Thu, Jan 30, 2025 5-52-07 PM.txt

R&D (FilterCavity)
Print this report.
RyutaroTakahashi - 17:17, Thursday 30 January 2025 (3898)Get code to link to this report
Replacement of DRY pump

I replaced the broken DRY pump (DSP500) with new one (NeoDry30G) for theTMP in the south end yesterday. The TMP is running with the DRY pump now.

Images attached to this report
3898_20250130091659_neodry30g.jpg
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 09:15, Wednesday 29 January 2025 (3895)Get code to link to this report
Comment to hydrogel measurement (Click here to view original report: 3890)

There was a use of positive sign convention at some places in the script instead of negative one. So, the plots of linear birefringence are updated as 

Non-image files attached to this comment
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 09:20, Thursday 23 January 2025 (3894)Get code to link to this report
Comment to hydrogel measurement (Click here to view original report: 3890)

I also reversed the plots to Stress vs. birefringence for better visualization.

Formula 1: Arylamide (There was a slight issue of this hydrogel slipping during measurement end)

Fig 1: Stress vs Retardation 

Fig 2: Stress vs. Diattenuation 

Formula 2: Di-phenyl

Fig 3: Stress vs Retardation 

Fig 4: Stress vs. Diattenuation 

Non-image files attached to this comment
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 16:07, Wednesday 22 January 2025 (3891)Get code to link to this report
Comment to hydrogel measurement (Click here to view original report: 3890)

The plots obtained after analysis for birefringence vs. stress.

Formula 1: Arylamide (There was a slight issue of this hydrogel slipping during measurement end)

Fig 1: Retardation vs. Stress

Fig 2: Diattenuation vs. Stress

Formula 2: Di-phenyl

Fig 3: Retardation vs. Stress

Fig 4: Diattenuation vs. Stress

Non-image files attached to this comment
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 11:58, Tuesday 21 January 2025 (3890)Get code to link to this report
hydrogel measurement

Katsuki, Shalika

We note down the value of change in width as well, from now on. This will impropve the stress estimation. The width mesaured is in last but one column and the force measured is in last column. The uncertainity on force is 0.01N, and  0.1mm on width and thickness. The measurement on width and thickness was done using vernier caliper. The thickness was the same all the time.

input_filename =  C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\Tue, Jan 21, 2025 9-55-00 AM.txt

Two formulas of hydrogel were measured.

hydrogel 1 (arylamide,1.5 mm, 25 mm )

output_filename = C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\Tue, Jan 21, 2025 10-18-53 AM.txt

hydrogel 2 (di-phenyl(YY037), 1.5 mm, 25 mm)

output_filename = C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\Tue, Jan 21, 2025 11-32-51 AM.txt

Comments related to this report
ShalikaSingh - 16:07, Wednesday 22 January 2025 (3891)

The plots obtained after analysis for birefringence vs. stress.

Formula 1: Arylamide (There was a slight issue of this hydrogel slipping during measurement end)

Fig 1: Retardation vs. Stress

Fig 2: Diattenuation vs. Stress

Formula 2: Di-phenyl

Fig 3: Retardation vs. Stress

Fig 4: Diattenuation vs. Stress

Non-image files attached to this comment
ShalikaSingh - 09:20, Thursday 23 January 2025 (3894)

I also reversed the plots to Stress vs. birefringence for better visualization.

Formula 1: Arylamide (There was a slight issue of this hydrogel slipping during measurement end)

Fig 1: Stress vs Retardation 

Fig 2: Stress vs. Diattenuation 

Formula 2: Di-phenyl

Fig 3: Stress vs Retardation 

Fig 4: Stress vs. Diattenuation 

Non-image files attached to this comment
ShalikaSingh - 09:15, Wednesday 29 January 2025 (3895)

There was a use of positive sign convention at some places in the script instead of negative one. So, the plots of linear birefringence are updated as 

Non-image files attached to this comment
BIGFOOT (General)
Print this report.
ShalikaSingh - 15:39, Monday 20 January 2025 (3888)Get code to link to this report
Comment to Labview modifications (Click here to view original report: 3886)

I have replaced the part which calculates the mean, with the statistics.vi to get median instead. 

BIGFOOT (General)
Print this report.
ShalikaSingh - 13:29, Monday 20 January 2025 (3886)Get code to link to this report
Labview modifications

In "Complete Integration- Dual Step.vi" I have added a place to add 2 extra columns, for some additional data saving as comments. Comments column will still be there if needed. We can save down the value of force and stress during measurement.

Comments related to this report
ShalikaSingh - 15:39, Monday 20 January 2025 (3888)

I have replaced the part which calculates the mean, with the statistics.vi to get median instead. 

KAGRA MIR (General)
Print this report.
MarcEisenmann - 18:34, Wednesday 15 January 2025 (3885)Get code to link to this report
new support for 30mm diameter samples

Previous 30mm diameter sample holder induced large stress due to mounting screw that was messing up the birefringence mapping.

After few iterations, the updated holder design seems reasonable and will be 3D printed tomorrow.

All the parts are saved in Dropbox and NAS folder 'mount and holder'

It should be possible to install all samples (ie with length 10,15,20,50,60,70 mm length).

Images attached to this report
3885_20250115103326_30mmholder.png
BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 22:00, Tuesday 14 January 2025 (3884)Get code to link to this report
Comment to Measurement to compare uncertainity (Click here to view original report: 3881)

0.5-3V with 0.15V step

input_filename = C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Tue, Jan 14, 2025 10-00-53 PM.txt

output_filename = C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Tue, Jan 14, 2025 11-31-23 AM.txt

BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 19:03, Monday 13 January 2025 (3881)Get code to link to this report
Measurement to compare uncertainity

Pol states for Sapphire measurements

0-3.5V with 0.1V step, 100 avg (previous setting)

input_filename  = C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Mon, Jan 13, 2025 12-44-19 PM.txt

output_filename  = C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Tue, Jan 14, 2025 8-04-23 AM.txt

0.5-3V with 0.25V step

input_filename = C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Mon, Jan 13, 2025 3-59-32 PM.txt

output_filename = C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Mon, Jan 13, 2025 4-34-12 PM.txt

Comments related to this report
ShalikaSingh - 22:00, Tuesday 14 January 2025 (3884)

0.5-3V with 0.15V step

input_filename = C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Tue, Jan 14, 2025 10-00-53 PM.txt

output_filename = C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Tue, Jan 14, 2025 11-31-23 AM.txt

BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 18:30, Sunday 12 January 2025 (3879)Get code to link to this report
Comment to new kind of hydrogel measurement (Click here to view original report: 3864)

The analysis of the results give the following results. 

Find the plot here: C:\Users\atama\Dropbox\Cell Birefringence\Measurement data\hydrogel\New bis-phenyl-1O 1.5 mm

Non-image files attached to this comment
BIGFOOT (Readout)
Print this report.
ShalikaSingh - 19:09, Friday 10 January 2025 (3878)Get code to link to this report
Trial to characterize EOM

On the eom path, made the laser linear using HWP and QWP (±0.01º). This was measured using polcam. The polcam was then removed. The power after this stage was 1.44mW.

Then aligned PD on the transmission path.

Then added polarizer infront of PD and the setup was configured to be in cross polarization configuration.

The characterization will be done using PD because the polcam can't see kHz modulation.

The test is being done to check how much modulation can be provided by the circuit.

BIGFOOT (Readout)
Print this report.
ShalikaSingh - 18:07, Friday 10 January 2025 (3873)Get code to link to this report
Beam requirement for EOM

The aperture of EOM is 2mm diameter. So, the beam radius should be smaller than 1mm.

The power density requirement is <4W/mm2.

The current beam characteristics are shown using lens F=100 in Fig 1 and F=50, -50 mm in Fig 2. from elog 3754

If we keep input power of around 1.5mW, both of the lens system would be fine to implement.

The 100 mm lens was kept at 101.5cm.

Tbh I didn't quite understand why the 100 mm lens was removed (in elog 3563). From the beam characteristics it seems we can achieve the requirement on both power density and beam size using one lens and in less space. I feel its much better to use less lens and keep stuff in compact space, unless there is some reasonable issue that needs to be taken into account. If the space is compact we can fit both EOM in reasonable space, without consuming the entire length of the table.

In Fig 3 I have added position for both EOM and the required lens for it.(will add the plot in a shortwhile)

Images attached to this report
3873_20250110063749_completebeambeforeandafterf100mm.png 3873_20250110063753_completebeambeforeandafter2f50mm.png
BIGFOOT (General)
Print this report.
ShalikaSingh - 17:28, Friday 10 January 2025 (3876)Get code to link to this report
Lightwave laser from tama input

The laser was tested to see if it can be switched on or is dead (the asset label says 1996). Apparently it's not dead. It's on bigfoot table now, and is intended to be used.

Images attached to this report
3876_20250110092614_20250110172628.jpg